OpenStudy (anonymous):

Write the equation of direct variation for x = -5 and y = 90

OpenStudy (mayankdevnani):

direct variation :- \[\huge \bf x=k \times y\] and \[\huge \bf k=\frac{x}{y}\] now, plug x=-5 and y=-90, \[\huge \bf k=\frac{-5}{-90}\] now what is the value of k?

OpenStudy (mayankdevnani):

@balokkk

OpenStudy (anonymous):

-0.6?

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

yeah. you are correct. @balokkk

OpenStudy (mayankdevnani):

@balokkk